OpenStudy (bossimbacon):

which value of b solves the equation b^-3= 1/216

OpenStudy (gorv):

b^-3=1/b^3 \[\frac{ 1 }{b ^{3}}=\frac{ 1 }{ 216 }=\frac{ 1 }{ 6^3 }\]

OpenStudy (gorv):

6^3=b^3 6=b

OpenStudy (bossimbacon):

tanks so much @gorv here is your medal and fan

OpenStudy (gorv):

welcome buddy :)